OpenStudy (anonymous):

3.14 x 10^6 is written how in scientific notation 314,000,000 ??

OpenStudy (anonymous):

it was 314,000,000 u r correct good job :)

OpenStudy (unklerhaukus):

not quite you have to move the Dot six places to the right so\[3.14 \times 10^6 = 3,140,000\]

OpenStudy (anonymous):

it was because you add 6 zeros

OpenStudy (anonymous):

nice job suzy4321

OpenStudy (unklerhaukus):

\[3 \times 10 ^6=3 000 000\] so\[ 3.14 \times 10^6=3 140 000\]

OpenStudy (anonymous):

so you add six after the decimal correct and if the exponent is negative it be to the left of the decimal right?

OpenStudy (unklerhaukus):

exactly , the decimal is the place holder and exponent on the ten will shift it